# The Orientation Boundary for Step-Duplicating Recursors: Mechanized Impossibility, Escape, and Certification

### 19 Mar 2026 v7

@@ 4.2 Affine / linear constructor-local measures
Every direct class blocked at the duplicating root step is also blocked for the full context-closed relation StepCtxFull, because every root contraction is a contextual contraction under the identity context. In particular, the additive, transparent-compositional, affine-with-pump, restricted-quadratic, bounded cross-term, bounded multilinear, max-plus, tracked-primary pair, weighted scalar-projection, and balanced mixed-coordinate families all admit explicit context-closed corollaries. Machine-checked in Meta/ContextClosedBarrier.lean as stepCtxFull_orientation_implies_root together with the corresponding class-specific KO7 corollaries.
This corollary does not add new obstruction mechanisms. Its role is semantic: it makes the root-step barrier stack available directly at the same context-closed level used by standard TRS tools and by the external TPDB / CeTA semantics in Section 6.
@@ 4.3 KO7 instantiation
Proof sketch. Direct instantiation of Theorem 4.1 with $\mathrm{base}=\mathrm{void}$, $\mathrm{succ}=\delta$, $\mathrm{wrap}=\mathrm{app}$, $\mathrm{recur}=\mathrm{rec}\Delta$. Machine-checked in Meta/CompositionalMeasure_Impossibility.lean. $\square$
@@ 4.4 Essentiality of the transparency hypothesis
$\mathrm{app}\;s\;(\mathrm{rec}\Delta\;b\;s\;n)\;<\;\mathrm{rec}\Delta\;b\;s\;(\delta\;n)$
@@
Once transparency is dropped, the Tier 2 class contains measures that orient the step. Machine-checked in Meta/RecCore.lean.
@@
Method Additive Transp.-comp. Affine Escape mechanism
$W$ (nonlinear poly.) nonlinear cross-variable coupling
KO7-specialized MPO recursive subterm comparison
DP projection modular SCC decomposition
@@ 4.5 The DP escape
Machine-checked in Meta/DP_BaseOrder_Boundary.lean as counterHeight_orients_dpPair, extracted_dp_problem_has_linear_base_order, and no_blanket_dp_poly_base_barrier. The point is boundary-setting rather than escape construction: modularity alone is not enough to force failure, but neither does it trivialize the direct barrier story. What matters is the transformed problem and the structure imported by the chosen base order.
@@ 4.6 Ablation: removing duplication dissolves the barrier
Machine-checked in Meta/LinearRec_Ablation.lean. Removing duplication reduces the problem to a simple additive size measure: the $\delta$-wrapper weight alone drives the decrease, with no need for the multiset or phase-bit components of $\mu^{3}_{c}$. This confirms that step-duplication is the operative source of every barrier theorem in this paper. Symmetric removal theorems for conditions (1) through (3) of §4 are not provided.
@@
This relaxation does not dissolve the direct barrier. The added collapse rule is harmless in isolation. It is strictly oriented by simpleSize, and the nonlinear full-step witness $W$ still proves root-step termination of the extended relation. However, the additive, transparent-compositional, and affine barrier theorems continue to apply because the duplicating rule rec_succ is unchanged. Machine-checked in Meta/ObjectAxiom_Ablation.lean as simpleSize_orients_app_collapse, wf_StepCollapseRev, no_global_stepCollapse_orientation_additive, no_global_stepCollapse_orientation_compositional_transparent, and no_global_stepCollapse_orientation_affine_of_unbounded. So this particular directed collapse surrogate is not the missing escape mechanism: it changes the system, but it does not neutralize the duplication obstruction isolated above. It is only a first-order syntactic proxy for collapse inside the rewrite language, not a replacement for full rewriting modulo an equational theory.
A separate artifact-only surrogate in Meta/SharingBarrierLift.lean makes the complementary point for condition (2) of §4: once the duplicated payload is represented through explicit sharing rather than copied in the tree, the tree-style direct barrier no longer applies verbatim. The module packages this as a positive counterexample (sharedCounter_orients_step) and a boundary theorem (sharing_breaks_tree_barrier), so the no-sharing hypothesis is not merely presentational.
Bounded SCC composite duplication.
Theorem 4.23 (Bounded delayed-duplication barrier).
@@
Each root rule already duplicates the payload once: the same term $s$ appears both as the left wrapper argument and again as the recursive step argument on the right-hand side. The point of the theorem is different: the two-node SCC yields a stable two-step composite profile and an induced minimal context relation on which the additive and affine barrier arguments still go through. Concretely, one root step followed by one step under the right argument of wrap realizes the effective pattern
@@
The concrete AltTerm development in Meta/MutualDuplication_Case.lean remains as the worked instance of this bounded SCC theorem. In that instance, the effective target carries three copies of the step payload measure against one copy on the source side, so the additive contradiction closes by the same coefficient comparison as in the schema barrier. The affine companion theorem uses the same two-step composite profile but packages it as a derived duplication schema with doubled successor and nested wrapper, so the existing affine barrier applies under the corresponding unbounded-pump hypothesis. The SCC layer now extends one step further: Meta/MutualDuplication_Transparent.lean proves that the same alternating composite is still blocked for transparent-compositional Tier 2 measures (no_global_orients_ctx_compositional_transparent) and for weighted functional matrix-style families whose projected scalar interface is already blocked (no_global_orients_ctx_matrixFunctional_of_projected_unbounded). A multiplicity-preserving SCC theorem still requires a different mechanism: in first-order tree rewriting, once every rule preserves the relevant payload multiplicities, the current coefficient-growth argument no longer has a duplicated payload to pump against.
@@
There exists a two-node first-order SCC together with an explicit payload-counting function such that each individual root rule preserves payload multiplicity exactly. On synchronized inputs carrying the same latent payload on two tagged channels, one full SCC cycle
@@
The source has no visible wrapper-carried payload, while the target exposes two visible wrapper copies of the synchronized payload and still preserves the counted multiplicity. Consequently no additive direct measure on this preserving SCC syntax can orient the synchronized composite uniformly. Equivalently, no such additive measure globally orients the induced minimal context relation generated by root steps and right-wrapper descent. Machine-checked in Meta/MutualDuplication_Preserving.lean as step_preserves_payloadCount, synchronized_cycle_realized, synchronized_cycle_exposes_payload, no_additive_orients_synchronized_cycle, and no_global_orients_ctx_additive.
@@
The mechanization also proves a conditional affine extension: under a wrapper-dominance condition $r_{p}+r_{q}<w_{l}+w_{r}\cdot w_{l}$ on the affine coefficients (where $r_{p},r_{q}$ are the recursor’s payload scaling coefficients and $w_{l},w_{r}$ are the wrapper’s left and right scaling coefficients), no affine measure with an unbounded pump can orient the synchronized cycle. The proof pumps the payload evaluation past the source constant, then exploits the coefficient gap to force the target above the source. Machine-checked as no_affine_orients_synchronized_cycle_of_wrapper_dominance in Meta/MutualDuplication_Preserving.lean.

### 08 Apr 2026 v8

@@ 4.2 Affine / linear constructor-local measures
Every direct class blocked at the duplicating root step is also blocked for the full context-closed relation StepCtxFull, because every root contraction is a contextual contraction under the identity context. In particular, the additive, transparent-compositional, affine-with-pump, restricted-quadratic, bounded cross-term, bounded multilinear, max-plus, tracked-primary pair, weighted scalar-projection, and balanced mixed-coordinate families all admit explicit context-closed corollaries. Proved in Meta/ContextClosedBarrier.lean as stepCtxFull_orientation_implies_root together with the corresponding class-specific KO7 corollaries.
This corollary adds no new obstruction mechanism. It simply states the root-step barrier stack at the same context-closed semantic level used by standard TRS tools and by the external TPDB / CeTA semantics in Section 6.
@@ 4.3 KO7 instantiation
Direct instantiation of Theorem 4.1 with $\mathrm{base}=\mathrm{void}$, $\mathrm{succ}=\delta$, $\mathrm{wrap}=\mathrm{app}$, $\mathrm{recur}=\mathrm{rec}\Delta$; see Meta/CompositionalMeasure_Impossibility.lean. $\square$
@@ 4.4 Essentiality of the transparency hypothesis
$CM.\mathrm{eval}(\mathrm{app}\;s\;(\mathrm{rec}\Delta\;b\;s\;n))\;<\;CM.\mathrm{eval}(\mathrm{rec}\Delta\;b\;s\;(\delta\;n))$
@@
Once transparency is dropped, the Tier 2 class contains measures that orient the step (see Meta/RecCore.lean).
@@
Table 5: Escape methods and the structural assumptions they import.
Escape method Broken / imported structural assumption Resulting escape shape
DP projection Wrapper sensitivity on the duplicated payload is discarded by projecting to the counter position. Modular transformed-relation escape.
Nonlinear $W$ witness Successor transparency and direct affine / additive structure are violated by cross-variable coupling. Global nonlinear full-step orientation.
KO7-specialized MPO The direct Nat-valued measure interface is left in favor of fixed-signature path-order structure, recursive subterm comparison, and a specialized same-head clause. Fixed-signature path-order escape.
@@ 4.5 The DP escape
The dependency-pair support layer has also been widened. Meta/DependencyPairs_FiniteGraph.lean reuses the finite-SCC search surface on any finite decidable dependency-pair relation and repackages a discovered SCC directly as the standard SCCCycle. Meta/DependencyPairs_CallGraph.lean and Meta/DependencyPairs_ExtractedCallGraph.lean remove the manual relation and finite-type packaging layers. Meta/DependencyPairs_TPDBExtraction.lean extracts the array-backed call graph from a finite TPDB-style rule list by computing defined heads and right-hand-side call heads, while Meta/DependencyPairs_FirstOrderExtraction.lean lifts the same step to a generic finite first-order TRS layer over arbitrary symbol and variable types.
The remaining frontend layers strip away packaging one stage at a time. Meta/DependencyPairs_FirstOrderFrontend.lean accepts an arbitrary finite rule-record array together with lhs / rhs extractors and reexports the same call-graph plus SCC-search / contradiction surface. Meta/DependencyPairs_FirstOrderEngine.lean packages that interface as a single internal engine object carrying Rule, rules, lhs, and rhs; it also bundles the typeclass-level first-order view and the raw-term bridge for engines that keep their own syntax. Meta/DependencyPairs_HeadView.lean works with the smaller head-view interface when only dependency-pair extraction matters. Meta/DependencyPairs_FiniteCarrierView.lean collects the finite rule-carrier, finite raw rule-carrier, finite head-carrier, and finite extracted-data carrier bridges in one module. Finally, Meta/DependencyPairs_KernelFirstOrder.lean packages the full KO7 root TRS directly over an internal symbol type and feeds that stack through the raw-term, head-view, and finite-carrier layers.
@@
The Lean proof is in Meta/DP_BaseOrder_Boundary.lean (counterHeight_orients_dpPair, extracted_dp_problem_has_linear_base_order, no_blanket_dp_poly_base_barrier). The point is boundary-setting rather than escape construction: modularity alone is not enough to force failure, but neither does it trivialize the direct barrier story. What matters is the transformed problem and the structure imported by the chosen base order.
@@ 4.6 Ablation: removing duplication dissolves the barrier
Formalized in Meta/LinearRec_Ablation.lean. Removing duplication reduces the problem to a simple additive size measure: the $\delta$-wrapper weight alone drives the decrease, with no need for the multiset or phase-bit components of $\mu^{3}_{c}$. This confirms that step-duplication is the operative source of the core immediate-step barrier theorems in this paper. Symmetric removal theorems for conditions (1) through (3) of §4 are not provided.
@@
This relaxation does not dissolve the direct barrier. The added collapse rule is harmless in isolation. It is strictly oriented by simpleSize, and the nonlinear full-step witness $W$ still proves root-step termination of the extended relation. However, the additive, transparent-compositional, and affine barrier theorems continue to apply because the duplicating rule rec_succ is unchanged. The Lean development is in Meta/ObjectAxiom_Ablation.lean; the relevant identifiers are simpleSize_orients_app_collapse, wf_StepCollapseRev, no_global_stepCollapse_orientation_additive, no_global_stepCollapse_orientation_compositional_transparent, and no_global_stepCollapse_orientation_affine_of_unbounded. So this particular directed collapse surrogate is not the missing escape mechanism: it changes the system, but it does not neutralize the duplication obstruction isolated above. It is only a first-order syntactic proxy for collapse inside the rewrite language, not a replacement for full rewriting modulo an equational theory.
A separate artifact-only surrogate in Meta/SharingBarrierLift.lean makes the complementary point for condition (2) of §4: once the duplicated payload is represented through explicit sharing rather than copied in the tree, the tree-style direct barrier no longer applies verbatim. The module packages this as a positive counterexample (sharedCounter_orients_step) and a boundary theorem (sharing_breaks_tree_barrier), so the no-sharing hypothesis is not merely presentational. Correspondingly, any higher-order lift that remains tree-like and forbids such sharing should still be expected to inherit the same duplication pressure; proving that conditional survival statement is outside the present first-order artifact.
Alternating-cycle composite-step barrier.
Theorem 4.24 (Alternating-cycle composite-step barrier).
@@
This theorem is not about non-duplicating local rules: each root rule already duplicates the payload once, since the same term $s$ appears both as the left wrapper argument and again as the recursive step argument on the right-hand side. Its content is that the two-node SCC yields a stable two-step composite profile and an induced minimal context relation on which the additive and affine barrier arguments still go through. Concretely, one root step followed by one step under the right argument of wrap passes through the intermediate trace $\mathrm{wrap}(s,\mathrm{recurB}(b,s,\mathrm{succ}(n)))$ and realizes the effective pattern
@@
The concrete AltTerm development in Meta/MutualDuplication_Case.lean remains the worked instance of this bounded SCC theorem. In that instance, the effective target carries three copies of the step payload measure against one copy on the source side, so the additive contradiction closes by the same coefficient comparison as in the schema barrier. The affine companion theorem uses the same two-step composite profile but repackages it as a derived duplication schema with doubled successor and nested wrapper, so the existing affine barrier applies under the corresponding unbounded-pump hypothesis. The delayed-duplication branch then extends to finite cyclic ‘k+1‘-node SCCs in Meta/MutualDuplication_KNode.lean. That file generalizes the alternating two-node construction to local rules $\mathrm{recur}_{i}(b,s,\mathrm{succ}(n))\to\mathrm{wrap}(s,\mathrm{recur}_{i+1}(b,s,n))$ modulo the cycle length, proves exact one-cycle realization (cycle_realized), and rederives both additive and affine minimal-context barriers for the finite cyclic family (no_global_orients_ctx_additive, no_global_orients_ctx_affine_of_unbounded).
That delayed-duplication finite-cycle branch is now factored one layer further through an abstract one-cycle interface in Meta/MutualDuplication_CycleFlow.lean. The abstract source state is $\mathrm{recur}(b,s,\mathrm{succ}^{m}(n))$, the abstract target state is $\mathrm{wrap}^{m}(s,\mathrm{recur}(b,s,n))$, and a certified cycle witness supplies the exact contextual path between them. On top of that witness, the file packages four generic contextual obstructions: additive (no_global_orients_ctx_additive), affine under the derived unbounded-pump hypothesis (no_global_orients_ctx_affine_of_unbounded), transparent-compositional with successor transparency at the base point (no_global_orients_ctx_compositional_transparent), and a scalar-projection lift reducing richer codomain orders to an affine contradiction on the derived one-cycle schema (no_global_orients_ctx_of_scalar_projection_affine_of_unbounded).
The companion bridge file Meta/MutualDuplication_KNode_Abstract.lean packages the finite cyclic ‘k+1‘-node delayed SCC as an instance of that abstract interface and rederives the additive, affine, transparent-compositional, and projected matrix-style corollaries as no_global_orients_ctx_additive_via_cycleFlow, no_global_orients_ctx_affine_of_unbounded_via_cycleFlow, no_global_orients_ctx_compositional_transparent_via_cycleFlow, and no_global_orients_ctx_matrixFunctional_of_projected_unbounded_via_cycleFlow. So the finite cyclic syntax is now the concrete witness, not the theorem shape itself. The earlier Meta/MutualDuplication_Transparent.lean file remains the original two-node transparent/projection presentation of the same delayed-duplication mechanism.
The delayed branch also has a raw-graph formulation in Meta/MutualDuplication_GraphCycle.lean. That module fixes an arbitrary node type and edge relation, equips the graph-indexed recursor family with a certified closed cycle witness (CyclePath, toCycleWitness), and proves that any such raw graph instance inherits the same additive, affine, transparent-compositional, and scalar-projection contextual barriers as the finite cyclic presentation (no_global_orients_ctx_additive, no_global_orients_ctx_affine_of_unbounded, no_global_orients_ctx_compositional_transparent, no_global_orients_ctx_of_scalar_projection_affine_of_unbounded).
The utility files Meta/GraphPathExtraction.lean and Meta/FiniteGraphReachability.lean remove most of the remaining witness plumbing. Any nonempty closed transitive-closure proof or round-trip SCC witness is converted automatically into the required raw-graph cycle via ofClosedTransGen and ofRoundTrip, and for finite decidable graphs the wrappers no_global_orients_ctx_additive_of_finiteRoundTrip and its affine / transparent / scalar-projection companions work directly from bounded reachability facts rather than explicit path proofs. Meta/FiniteGraphSCC.lean then adds a finite search surface on node pairs: a candidate-pair finset sccPairs, an optional SCC-pair search function, and an equivalence between successful search and HasNontrivialSCC. From there the wrappers no_global_orients_ctx_additive_of_hasNontrivialSCC and its affine / transparent / scalar-projection companions recover the concrete witness pair and both reachability directions internally.
The relation-level wrapper file Meta/MutualDuplication_RelationalGraph.lean removes one more manual layer. Starting only from a node type, edge relation, shared constructor interface, and local edge-realization theorem, it builds the delayed raw-graph system automatically and reexports the same round-trip / finite-round-trip / HasNontrivialSCC wrappers together with direct variants that consume a successful finite-SCC search result. Two concrete frontends sit on top of that layer. Meta/MutualDuplication_CallGraph.lean starts from extracted call-graph data (‘nodeKey‘, ‘succKeys‘) on a finite node type and builds the same delayed SCC interface automatically, while Meta/MutualDuplication_ExtractedCallGraph.lean removes the finite-type packaging too: raw extracted node data stored in an array yields the node type ‘Fin nodes.size‘, the call graph, the finite SCC search surface, and then the delayed barrier wrappers. The delayed SCC obstruction is therefore packaged at ten levels: explicit finite cycle, abstract one-cycle witness, raw graph with a certified cycle object, raw graph with automatic cycle extraction from a closed-path witness, raw graph with finite-round-trip reachability wrappers, raw graph with finite-SCC witness packaging, raw graph with finite search-result wrappers, relation-level construction of the raw graph, call-graph construction from extracted successor-key data, and array-backed construction from raw extracted node data.
A multiplicity-preserving SCC theorem still requires a different mechanism: in first-order tree rewriting, once every rule preserves the relevant payload multiplicities, the current coefficient-growth argument no longer has a duplicated payload to pump against. Delayed branch: per-rule duplication, composite-step barrier. Preserving branch: per-rule multiplicity preserved, synchronized-exposure barrier.
@@
There exists a two-node first-order SCC together with an explicit payload-counting function such that each individual root rule preserves payload multiplicity exactly. On synchronized inputs carrying the same latent payload on two tagged channels, one full SCC cycle realizes the contextual reduction
@@
The source has no visible wrapper-carried payload, while the target exposes two visible wrapper copies of the synchronized payload and still preserves the counted multiplicity. Consequently no additive direct measure on this preserving SCC syntax can orient the synchronized composite uniformly. Equivalently, no such additive measure globally orients the induced minimal context relation generated by root steps and right-wrapper descent. Formalized in Meta/MutualDuplication_Preserving.lean as step_preserves_payloadCount, synchronized_cycle_realized, synchronized_cycle_exposes_payload, no_additive_orients_synchronized_cycle, and no_global_orients_ctx_additive.
@@
The mechanization also proves a conditional affine extension: under a wrapper-dominance condition $r_{p}+r_{q}<w_{l}+w_{r}\cdot w_{l}$ on the affine coefficients (where $r_{p},r_{q}$ are the recursor’s payload scaling coefficients and $w_{l},w_{r}$ are the wrapper’s left and right scaling coefficients), no affine measure with an unbounded pump can orient the synchronized cycle. The proof pumps the payload evaluation past the source constant, then exploits the coefficient gap to force the target above the source. Proved as no_affine_orients_synchronized_cycle_of_wrapper_dominance in Meta/MutualDuplication_Preserving.lean.
A finite cyclic synchronized-packet extension appears in Meta/MutualDuplication_Preserving_KNode.lean. There the latent payload is carried by an explicit packet of indexed channels of length ‘k+1‘; one root step exposes the current packet head under wrap and forwards the remaining packet to the next cycle node. The file proves exact payload preservation for every local step (step_preserves_payloadCount), exact one-cycle realization (synchronized_cycle_realized), additive failure of the induced minimal context relation (no_global_orients_ctx_additive), and a conditional affine failure theorem under a generalized wrapper-dominance condition (no_global_orients_ctx_affine_of_wrapper_dominance). This is a narrowed explicit payload-flow model rather than a fully abstract SCC metatheorem, but it shows that the synchronized-exposure obstruction survives beyond the original two-node presentation.
That explicit packet model is then factored through an abstract payload-flow interface in Meta/MutualDuplication_PayloadFlow.lean. The companion bridge file Meta/MutualDuplication_Preserving_Abstract.lean packages the finite synchronized-packet SCC as an instance of that interface and rederives the additive and affine global barrier theorems from the abstract cycle lemmas no_global_orients_ctx_additive and no_global_orients_ctx_affine_of_wrapper_dominance. It also rederives the packet-transparent and projected packet-transparent corollaries via no_global_orients_ctx_packetTransparent_via_payloadFlow and no_global_orients_ctx_matrixFunctional_of_projected_packetTransparent_via_payloadFlow. The packet syntax is therefore a concrete witness presentation, not the theorem shape itself: the barrier mechanism has been isolated as a generic one-cycle payload-flow metatheorem.
A narrow transparent-compositional downstream corollary is also available in Meta/MutualDuplication_Preserving_Transparent.lean. That module treats the synchronized packet constructors as evaluation-transparent bookkeeping and proves both the cycle-level contradiction no_compositional_orients_synchronized_cycle_packetTransparent and the induced minimal-context corollary no_global_orients_ctx_packetTransparent. It also adds a generic scalar-projection lift, no_global_orients_ctx_matrixFunctional_of_projected_packetTransparent: any componentwise matrix-style orienter whose chosen weighted scalar projection collapses to one of these packet-transparent measures is blocked as well. The preserving finite-cycle branch thus admits both a direct transparent obstruction and a projection-based downstream corollary, without introducing a separate preserving-specific matrix syntax layer.
The preserving branch also has a raw-graph formulation in Meta/MutualDuplication_PacketGraph.lean. That module fixes an arbitrary node type and edge relation, equips the graph-indexed packet system with a certified closed cycle witness (CyclePath, toCycleWitness), and proves additive, affine, transparent, and scalar-projection contextual barriers directly for that raw graph presentation (no_global_orients_ctx_additive, no_global_orients_ctx_affine_of_wrapper_dominance, no_global_orients_ctx_transparent, no_global_orients_ctx_of_scalar_projection_transparent).
Here too the path-extraction and finite-reachability utilities remove most of the cycle bookkeeping. A nonempty closed transitive-closure proof or round-trip SCC witness builds the required cycle object automatically via ofClosedTransGen and ofRoundTrip; for finite decidable graphs the *_of_finiteRoundTrip wrappers apply directly from bounded reachability facts. The finite-SCC layer in Meta/FiniteGraphSCC.lean goes one step further: it exposes a finite pair-search surface built around sccPairs, proves that successful search is equivalent to HasNontrivialSCC, and lets the wrappers no_global_orients_ctx_additive_of_hasNontrivialSCC and its affine / transparent / scalar-projection companions recover the concrete pair and both reachability directions internally.
The same relation-level wrapper file, Meta/MutualDuplication_RelationalGraph.lean, builds the preserving raw-graph system from the smaller relation-level presentation and reexports the same round-trip / finite-round-trip / HasNontrivialSCC wrappers together with direct variants that consume a successful finite-SCC search result. On top of that, Meta/MutualDuplication_CallGraph.lean and Meta/MutualDuplication_ExtractedCallGraph.lean provide the same preserving barrier surface starting respectively from extracted call-graph data on a finite node type and from raw extracted node data stored in an array. The obstruction is therefore no longer tied to finite packet syntax, hand-written cycle indices or explicit transitive-closure witnesses, manual raw-graph packaging, hand-supplied SCC pairs or separate SCC-existence certificates once finite search succeeds, or a hand-built finite node type / call-graph layer for extracted dependency data.